# Theoretical Investigation on Inductive Bias of Isolation Forest

### v1

@@ Appendix A Proofs in Section 3
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i+1},\mathbf{\Theta}\right)\right]$ $=\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]+1$
@@
Assuming that Eq. (2) holds for all $n<k\ (k>i+1)$, we then consider $n=k$. There are three cases: If the first split occurs in $[x^{\prime},x_{1}]$, the expected depth of $x_{i}$ increases by 1; If the first split occurs in $[x_{1},x_{i}]$, the expected depth of $x_{i}$ remains unchanged; otherwise, a split in $[x^{\prime},x_{i}]$ will eventually occur, and the expected depth of $x_{i}$ increases by about $\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}$. Formally, we have
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{k},\mathbf{\Theta}\right)\right]$ $=\frac{x_{1}-x^{\prime}}{x_{k}-x^{\prime}}(1+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{k},\mathbf{\Theta}\right)\right])$
@@
$\quad\quad+\sum_{j=i+1}^{k}\frac{x_{j}-x_{j-1}}{x_{k}-x^{\prime}}\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{j-1},\mathbf{\Theta}_{2:\tau}\right)\right]\ .$
By mathematical induction, we have
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{j-1},\mathbf{\Theta}_{2:\tau}\right)\right]=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{j-1},\mathbf{\Theta}_{2:\tau}\right)\right]\ ,$
@@
$\sum_{j=i+1}^{k}\frac{x_{j}-x_{j-1}}{x_{k}-x^{\prime}}\mathbb{E}_{\mathbf{\Theta}}$ $\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{j-1},\mathbf{\Theta}_{2:\tau}\right)\right]$
@@
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{k},\mathbf{\Theta}\right)\right]=$ $\ \frac{x_{1}-x^{\prime}}{x_{k}-x^{\prime}}+\frac{x_{k}-x_{i}}{x_{k}-x^{\prime}}\cdot\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+$
@@
A.5 Depth Function of Points Outside the Range of the Dataset
For data points outside the range of the dataset, we have the following theorem.
@@ A.5 Depth Function of Points Outside the Range of the Dataset
Proof   For $x<x_{1}$ or $x>x_{n}$, evidently, isolating $x$ is equivalent to isolating $x_{1}$ or $x_{n}$, respectively. Thus, we have
@@
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x;x_{1},\dots,x_{n},\mathbf{\Theta}\right)\right]=\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{n};x_{1},\dots,x_{n},\mathbf{\Theta}\right)\right],\textrm{for }\forall x>x_{n}\ .$
We now consider the case where $x\in\left[x_{1},x_{n}\right]$. By Lemma 3.4, it suffices to show that for $x\in\left[x_{i},x_{i+1}\right]$, the depth function $h\left(x;x_{1},\dots,x_{n},\mathbf{\Theta}\right)$ is a linear function. For any given $\mathbf{\Theta}=\mathbf{\Theta}_{0}$, we have
@@
$\frac{\partial L}{\partial s_{i}}=&\ \frac{U_{\textrm{ms}}\cdot\kappa+s_{2}+\dots+s_{i-1}}{(U_{\textrm{ms}}\cdot\kappa+s_{2}+\dots+s_{i-1}+s_{i})^{2}}-\sum_{i^{\prime}=i+1}^{n-2}{\frac{s_{i^{\prime}}}{(U_{\textrm{ms}}\cdot\kappa+s_{2}+\dots+s_{i^{\prime}})^{2}}}\\ &\ -\frac{U_{\textrm{ms}}}{U_{\textrm{ms}}\cdot\kappa+s_{2}+\dots+s_{n-2}+U_{\textrm{ms}}}-\mu_{i}+\mu_{i}^{\prime}\\ =&\ 0\ ,\end{aligned}$
@@
$U_{\textrm{cs}}=\max_{j\neq n_{0}/2}\lvert x_{j+1}-x_{j}\rvert\quad L_{\textrm{cs}}=\min_{j\neq n_{0}/2}\lvert x_{j+1}-x_{j}\rvert\ .$

### v2

@@ Appendix A Proofs in Section 3
$\mathbb{E}_{\mathbf{\Theta}^{\prime}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i+1},\mathbf{\Theta}^{\prime}\right)\right]$ $=\mathbb{E}_{\mathbf{\Theta}^{\prime}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}^{\prime}\right)\right]+1$
@@
Assuming that Eq. (2) holds for all n<k(k>i+1)n<k\ (k>i+1), we then consider $n=k$. There are three cases: If the first split occurs in $x^{\prime},x_{1}]$, the expected depth of $x_{i}$ increases by 1; If the first split occurs in $x_{1},x_{i}]$, the expected depth of $x_{i}$ remains unchanged; otherwise, a split in $x^{\prime},x_{i}]$ will eventually occur, and the expected depth of $x_{i}$ increases by about $\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}$. Formally, we have
$\mathbb{E}_{\mathbf{\Theta}^{\prime}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{k},\mathbf{\Theta}^{\prime}\right)\right]$ $=\frac{x_{1}-x^{\prime}}{x_{k}-x^{\prime}}(1+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{k},\mathbf{\Theta}\right)\right])$
@@
$\quad\quad+\sum_{j=i+1}^{k}\frac{x_{j}-x_{j-1}}{x_{k}-x^{\prime}}\mathbb{E}_{\mathbf{\Theta}^{\prime}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{j-1},\mathbf{\Theta}^{\prime}\right)\right]\ .$
By induction hypothesis, we have
$\mathbb{E}_{\mathbf{\Theta}^{\prime}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{j-1},\mathbf{\Theta}^{\prime}\right)\right]=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{j-1},\mathbf{\Theta}\right)\right]\ ,$
@@
$\sum_{j=i+1}^{k}\frac{x_{j}-x_{j-1}}{x_{k}-x^{\prime}}\mathbb{E}_{\mathbf{\Theta}^{\prime}}$ $\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{j-1},\mathbf{\Theta}^{\prime}\right)\right]$
@@
$\mathbb{E}_{\mathbf{\Theta}^{\prime}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{k},\mathbf{\Theta}^{\prime}\right)\right]=$ $\ \frac{x_{1}-x^{\prime}}{x_{k}-x^{\prime}}+\frac{x_{k}-x_{i}}{x_{k}-x^{\prime}}\cdot\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+$
@@
A.5 Depth Function of Points Not in the Dataset
For data points not in the dataset, we have the following theorem.
@@ A.5 Depth Function of Points Not in the Dataset
Proof For $x<x_{1}$ or $x\geq x_{n}$, evidently, isolating $x$ is equivalent to isolating $x_{1}$ or $x_{n}$, respectively. Thus, we have
@@
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x;x_{1},\dots,x_{n},\mathbf{\Theta}\right)\right]=\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{n};x_{1},\dots,x_{n},\mathbf{\Theta}\right)\right],\textrm{for }\forall x\geq x_{n}\ .$
We now consider the case where $x\in\left[x_{1},x_{n}\right)$. By Lemma 3.4, it suffices to show that for $x\in\left[x_{i},x_{i+1}\right)$, the expected depth $\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x;x_{1},\dots,x_{n},\mathbf{\Theta}\right)\right]$ is a linear function. For any given $\mathbf{\Theta}=\mathbf{\Theta}_{0}$, we have
@@
$\begin{aligned} \frac{\partial L}{\partial s_{i}}=&\ \frac{U_{\textrm{ms}}\cdot\kappa+s_{2}+\dots+s_{i-1}}{(U_{\textrm{ms}}\cdot\kappa+s_{2}+\dots+s_{i-1}+s_{i})^{2}}-\sum_{i^{\prime}=i+1}^{n-2}{\frac{s_{i^{\prime}}}{(U_{\textrm{ms}}\cdot\kappa+s_{2}+\dots+s_{i^{\prime}})^{2}}}\\ &\ -\frac{U_{\textrm{ms}}}{U_{\textrm{ms}}\cdot\kappa+s_{2}+\dots+s_{n-2}+U_{\textrm{ms}}}-\mu_{i}+\mu_{i}^{\prime}\\ =&\ 0\ ,\end{aligned}$
@@
$U_{\textrm{cs}}=\max_{j\neq n_{0}/2}\lvert x_{j+1}-x_{j}\rvert\quad L_{\textrm{cs}}=\min_{j\neq n_{0}/2}\lvert x_{j+1}-x_{j}\rvert\$.